Bartolo Delivery has two divisions, air express and ground service, that share the common costs of the company’s communications network, which are $8,000,000 a year. You have the following information about the two divisions and the common communications network:

Calls (thousands) Time on Network (hours)
Air express 574,000 350,000
Ground service 246,000 1,050,000

Required:

Determine the cost allocation if $5.60 million of the communications network costs are fixed and allocated on the basis of time on network, and the remaining costs, which are variable, are allocated on the basis of the number of calls. (Do not round intermediate calculations. Enter your answers in thousands of dollars.)
Answer #1
Call Time of network
Air express 574,000 350,000
Ground service 246,000 1,050,000
air expenses
Fixed =5600000*350000/(350000+1050000) 1400000
Variable =2400000*574000/(574000+246000) 1680000
Ground service
Fixed =5600000*1050000/(350000+1050000) 4200000
Variable =2400000*246000/(574000+246000) 720000
Therefore final solution
air expenses Ground service
Fixed 1400000 4200000
Variable 1680000 720000
Total 3080000 4920000